Find the dimensions of a rectangle whose length is 6 inches greater than its width. The area is 72 square inches. Show your work and explain the steps you used to find the solution in order to receive full credit.
Answer by DENNIS DON(8)

let L and W be the length and width of the rectangle
L=W+6
Area=L*W
= (W+6)*W
= W²+6W
also;
area=72
then;
W²+6W=72
W²+6W-72=0
(W-6)(W+12)=0
W=6 or W=-12
BUT the dimension can't be negative
then;
W=6, and
L=6+6
=12